OpenStudy (anonymous):

Use the elimination method to solve the following system of equations. x + 3y – z = 2 x – 2y + 3z = 7 x + 2y – 5z = –21 (2, 3, 5) (–2, 3, 5) (2, –3, 5) (2, 3, –5)

OpenStudy (phi):

elimination is a methodical process. Often people start by writing just the coefficients (but keeping all the variables "lined up" in alphabetical order like ou have...

OpenStudy (phi):

for example, start with 1 3 -1 2 1 -2 3 7 1 2 -5 -21 you start by zeroing out the numbers below 1 in the first column to zero out the 1 in the 2nd row, multiply the first row by -1 and add that whole row to the 2nd row. You keep the first row as it is

OpenStudy (phi):

1 3 -1 2 mutiply by -1 to get -1 -3 1 -2 add that to the 2nd row 1 -2 3 7 to get 0 -5 4 5 that is your new 2nd row 1 3 -1 2 0 -5 4 5 1 2 -5 -21

OpenStudy (phi):

can you zero out the 1 in the 3rd row ? In this case all you do is add -1 -3 1 -2 to the 3rd row

OpenStudy (anonymous):

start with the first two equations and eliminate a variable...I eliminated x x + 3y - z = 2 -->(-1)x + 3y - z = 2 x - 2y + 3z = 7 --------------- -x - 3y + z = - 2 (result of multiplying by -1) x - 2y + 3z = 7 ---------------add 0 - 5y + 4z = 5 -5y + 4z = 5 now take the last two equations and eliminate x x - 2y + 3z = 7 -->(-1)x - 2y + 3z = 7 x + 2y - 5z = -21 ---------------- -x + 2y - 3z = -7 (result of multiplying by -1) x + 2y - 5z = - 21 -----------------add 0 + 4y - 8z = - 28 4y - 8z = - 28 now take the answers you got...the ones where the x's are eliminated -5y + 4z = 5 -->(2)-5y + 4z = 5 4y - 8z = - 28 ----------------- -10y + 8z = 10 (result of multiplying by 2) 4y - 8z = - 28 ----------------add -6y = -18 y = - 18/-6 y = 3 now sub 3 in for y in either of the last two equations... 4y - 8z = - 28 4(3) - 8z = - 28 12 - 8z = - 28 -8z = - 28 - 12 -8z = - 40 z = -40/-8 z = 5 now sub 5 in for z and 3 in for y in any of the original equations... x + 3y - z = 2 x + 3(3) - 5 = 2 x + 9 - 5 = 2 x + 4 = 2 x = 2 - 4 x = - 2 check... x - 2y + 3z = 7 -2 - 2(3) + 3(5) = 7 -2 - 6 + 15 = 7 -8 + 15 = 7 7 = 7 (correct) ANSWER : (-2, 3, 5) Is there anything you do not understand ? If there is please ask. I will be happy to explain any of the work I have done. It is not hard, just time consuming.